# Continuous K-space Recovery Network with Image Guidance for Fast MRI Reconstruction

### v1

@@ III-A1 MRI Reconstrucion
III-A2 Implicit Neural Representation
In contrast to the typical representation of images using discrete pixel arrays, implicit neural representation (INR) models the mapping relationship between image coordinates $C$ and their corresponding intensity values $i$ as a function, which can be formulated as:
$i=f_{\theta}(C,V_{c})$ (3)
Here, $V_{c}$ represents the feature vector specific to the pixel at spatial coordinates $C$ and is obtained by establishing a latent space by training an encoder. Function $f_{\theta}$ is fitted using a multilayer perceptron (MLP) and is continuous to the spatial coordinate system. Therefore, once $f_{\theta}$ is well-fitted, intensity values can be queried at arbitrary coordinates, thereby achieving a continuous visual representation of the image.
III-A3 continuous K-space Recovery Network via INR
As described in (1), it is common to undersample the k-space for accelerated MRI. The correct recovery of the unsampled k-values, i.e., $K_{us}$, is crucial for high-quality MRI reconstruction. Drawing inspiration from the continuous image modeling of INR, we customize a k-space recovery net. Specifically, $K_{us}$ can be queried by using their coordinates $C_{us}$:
$K_{{us}}=f_{\theta}(C_{{us}},V_{C_{us}})$ (4)
$K_{{us}}=\Phi_{Dec}(C_{{us}},\Phi_{Enc}(K_{s},C_{s}))$ (5)
Here, the encoder $\Phi_{Enc}$ and decoder $\Phi_{Dec}$ are implemented through the transformer attention mechanism.
III-B Network Architecture
As shown in Fig.2, the overall architecture of our IGKR-Net can be roughly divided into two parts, i.e., the primary part directly processing in k-space domain (denoted as red lines), and the auxiliary part operating on the data of image domain (denoted as blue lines). The k-space part comprises two modules, i.e., LRIT and HRIT, to reconstruct the k-space from low resolution (LR) to high resolution (HR) progressively, while the image part comprises IDGM and TRAM for image-domain guidance and final refinement, respectively.
Specifically, given undersampled k-space $K\in R^{2\times H\times W}$ as input, we obtain the set of k-values $K_{s}$ and the corresponding coordinates $C_{s}$, from the sampled points. Then, $K_{s}$ and $C_{s}$ are encoded separately, and the resulting features are summed to form a vector sequence which is used as the input of LRIT. LRIT first uses a transformer-based encoder to learn a latent space and then recover a complete while LR k-space $\widehat{{K}}_{1}$ by querying pre-designed LR coordinates through the decoder. Subsequently, we get $\widehat{{I}}_{1}$ through IFFT and send it into IDGM to deeply mine semantic information from the corrupted image $I_{s}$ and guide the k-space recovery. We fuse $I_{s}$ with $\widehat{{I}}_{1}$, obtaining $\widehat{{I}}_{2}$ and $\widehat{{K}}_{2}$. Next, HRIT uses the recovered dense $\widehat{{K}}_{2}$ to further establish a continuous latent space, recovering complete and HR $\widehat{K}_{3}$ and its corresponding $\widehat{{I}}_{3}$ using the HR coordinates. Finally, we design TARM to refine $\widehat{{I}}_{3}$ in the image domain, resulting in the final output, i.e., $\widehat{{I}}_{4}$ and $\widehat{{K}}_{4}$.
@@ III-B Network Architecture
III-B1 Low Resolution Implicit Transformer (LRIT)
@@
LRIT and HRIT share the same structure but differ in their inputs. Here, we take LRIT as an example to explain its structure in detail. As depicted in Fig.2 (a), LRIT is an encoder-decoder architecture aiming to recover LR k-space $\widehat{{K}}_{1}$ and its corresponding $\widehat{{I}}_{1}$ by querying the LR coordinates.
@@
Specifically, given a set of sampled k-values $K_{s}$ and its coordinates $C_{s}$, a tokenization procedure is firstly applied to convert them into a vector sequence $\mathcal{V}$:
$\mathcal{V}=MLP(K_{s})+PE(C_{s})$ (6)
@@
where a multilayer perceptrons (MLPs) layer is applied to $K_{s}$, and PE refers to positional encodings with sine and cosine functions. Then we send $\mathcal{V}$ into the encoder of LRIT which consists of multi-head self-attention (MSA) and feed-forward network (FFN) to obtain the latent space $\mathcal{S}$ as follows:
$\mathcal{S}=FFN(MSA(\mathcal{V}))$ (7)
@@
Subsequently, we design a decoder that employs standard transformer decoder layers consisting of multi-head cross attention (MCA), MSA, and FFN, to recover complete while low-resolution k-space, i.e., $\widehat{{K}}_{1}$. In detail, for MCA, we employ encoded LR coordinates ${C}_{lr}\in R^{\frac{H}{2}\times\frac{W}{2}}$ to serve as the query ($Q$) and apply two different linear projects to latent space $\mathcal{S}$ to serve as the key ($K$) and value ($V$). After that, we apply MSA to the weighted feature again and decode it via FFN. Thus, we obtain a low-resolution while complete k-space $\widehat{{K}}_{1}\in R^{2\times\frac{H}{2}\times\frac{W}{2}}$, as well as its corresponding MRI image $\widehat{{I}}_{1}\in R^{1\times\frac{H}{2}\times\frac{W}{2}}$. Which can be formulated as:
$\widehat{{K}}_{1}=FFN(MSA(MCA(Q,K,V))\vspace{-0.05cm}$ (8)
$\widehat{{I}}_{1}={IFFT}(\widehat{{K}}_{1})$ (9)
III-B2 Image Domain Guidance Module (IDGM)
As shown in Fig.2 (b), our IDGM fuses the low-quality image $I_{s}$ and the output of LRIT, i.e., $\widehat{{I}}_{1}$, to fully utilize the original image information and guide the process of k-space recovery.
Fig.3 (a) illustrates the detailed structure of the IDGM, which can be divided into two stages: shallow stage and deep stage. Specifically, during the shallow stage, we firstly perform two convolutions with different strides to extract shallow features from $I_{s}$ and $\widehat{{I}}_{1}$ and unify them to the same size, i.e., ${\widehat{I}_{1}}\in R^{1\times\frac{H}{2}\times\frac{W}{2}}\to R^{h\times\frac{H}{2}\times\frac{W}{2}}$, ${I_{s}}\in R^{1\times H\times W}\to R^{h\times\frac{H}{2}\times\frac{W}{2}}$, where h denotes the channel number. Subsequently, the features are fused via element-wise summation and restored to ${{I}}_{up}\in R^{1\times H\times W}$ via another convolution and upsampling $up(\cdot)$, i.e., $I_{up}=up(conv(conv({\widehat{{I}}_{1}})+conv(I_{s})))$.
During the deep fusion, we first employ two convolution heads to extract image-specific shallow features from $I_{s}$ and ${{I}}_{up}$, respectively. After element-wise summation, the fused feature is sent into two attention branches, generating local channel attention $A_{l}$ without any pooling and global channel attention $A_{g}$ by using global average pooling, respectively. Given $A_{l}$ and $A_{g}$, the final attention weights $A$ can be obtained by the broadcasted addition (+), i.e., $A=sigmoid(A_{l}\textbf{+}A_{g})$. Finally, we obtain $\widehat{{I}}_{2}$, as well as its corresponding k-space $\widehat{{K}}_{2}$:
$\widehat{{I}}_{2}=conv(A\otimes conv({I}_{up}))\vspace{-0.06cm}$ (10)
$\widehat{{K}}_{2}={FFT}(\widehat{{I}}_{2})$ (11)
III-B3 High Resolution Implicit Transformer (HRIT)
As shown in Fig.2 (c), to further reconstruct the dense k-space with high resolution, we apply the implicit transformer again. Different from the above LRIT, here we use the $K_{s}^{\prime}$ and $C_{s}^{\prime}$ from restored $\widehat{{K}}_{2}$ to form vector sequence $\mathcal{V^{\prime}}$, i.e., $\mathcal{V^{\prime}}=MLP(K_{s}^{\prime})+PE(C_{s}^{\prime})$. Then we get latent space $\mathcal{S^{\prime}}$:
$\mathcal{S^{\prime}}=FFN(MSA(\mathcal{V^{\prime}}))$ (12)
Finally, we obtain the high quality k-space $\widehat{{K}}_{3}$ and its corresponding MRI image $\widehat{{I}}_{3}$ by encoding HR coordinates $C_{hr}\in R^{H\times W}$ as query ($Q^{\prime}$) and projecting $\mathcal{S^{\prime}}$ as key ($K^{\prime}$) and value ($V^{\prime}$). Which can be formulated as:
$\widehat{{K}}_{3}=FFN(MSA(MCA(Q^{\prime},K^{\prime},V^{\prime}))\vspace{-0.06cm}$ (13)
$\widehat{{I}}_{3}={IFFT}(\widehat{{K}}_{3})$ (14)
TABLE I: Quantitative results of the comparison experiments on the CC359 validation dataset using different masks. A higher value of metric↑ indicates better performance, whereas a lower value of metric↓ corresponds to improved performance.
Method PSNR↑ SSIM↑ NMSE↓ LPIPS↓ PSNR↑ SSIM↑ NMSE↓ LPIPS↓
ResNet[56] 28.13±3.84 0.8559±0.14 0.0200±0.01 0.0858±0.07 30.05±3.58 0.9052±0.07 0.0131±0.01 0.0569±0.04
UNet[26] 29.34±4.59 0.8756±0.11 0.0156±0.01 0.0618±0.05 31.73±4.70 0.9189±0.07 0.0090±0.00 0.0408±0.04
SwinMR[36] 31.11±3.87 0.9006±0.09 0.0124±0.00 0.0413±0.03 35.67±4.13 0.9543±0.04 0.0045±0.00 0.0184±0.02
RefineGAN[57] 29.96±4.59 0.8888±0.07 0.0303±0.02 0.0395±0.04 38.57±5.16 0.9656±0.02 0.0058±0.01 0.0077±0.01
SwinGAN[58] 31.32±3.20 0.9017±0.06 0.0087±0.01 0.0390±0.02 38.67±4.56 0.9688±0.02 0.0043±0.01 0.0156±0.02
OCUCFormer[59] 32.54±2.56 0.9166±0.06 0.0074±0.02 0.0377±0.03 38.70±5.33 0.9690±0.03 0.0035±0.02 0.0150±0.01
ResNet[56] 28.83±3.47 0.8576±0.12 0.0185±0.02 0.0841±0.08 31.00±3.66 0.9102±0.09 0.0181±0.05 0.0592±0.10
UNet[26] 30.39±4.78 0.8790±0.12 0.0124±0.00 0.0578±0.05 33.39±4.98 0.9280±0.08 0.0063±0.00 0.0362±0.03
SwinMR[36] 31.44±4.01 0.8943±0.09 0.0116±0.04 0.0441±0.03 34.36±3.85 0.9378±0.05 0.0059±0.00 0.0220±0.01
RefineGAN[57] 30.08±4.61 0.8770±0.07 0.0303±0.02 0.0395±0.03 35.73±5.05 0.9486±0.03 0.0098±0.01 0.0140±0.01
SwinGAN[58] 34.56±3.56 0.9233±0.04 0.0078±0.01 0.0221±0.02 38.89±4.32 0.9676±0.04 0.0046±0.02 0.0089±0.01
OCUCFormer[59] 35.55±3.20 0.9320±0.03 0.0050±0.02 0.0199±0.01 40.41±3.89 0.9770±0.04 0.0027±0.01 0.0059±0.01
TABLE II: Quantitative results of the comparison experiments on the fastMRI validation dataset using different masks. A higher value of metric↑ indicates better performance, whereas a lower value of metric↓ corresponds to improved performance.
Method PSNR↑ SSIM↑ NMSE↓ LPIPS↓ PSNR↑ SSIM↑ NMSE↓ LPIPS↓
ResNet[56] 24.54±2.02 0.6062±0.09 0.0461±0.02 0.2469±0.08 25.75±2.04 0.7059±0.07 0.0351±0.02 0.1791±0.06
UNet[26] 26.53±2.22 0.6599±0.10 0.0316±0.02 0.2242±0.08 28.21±2.33 0.7559±0.07 0.0220±0.02 0.1541±0.06
SwinMR[36] 24.67±2.21 0.6214±0.10 0.0480±0.03 0.2166±0.06 27.33±2.46 0.7530±0.08 0.0271±0.02 0.1354±0.05
RefineGAN[57] 25.71±2.56 0.6961±0.06 0.0335±0.02 0.1922±0.04 29.38±2.37 0.8054±0.02 0.0174±0.08 0.1376±0.01
SwinGAN[58] 26.60±2.09 0.7001±0.08 0.0330±0.01 0.1996±0.03 29.89±2.33 0.8069±0.06 0.0169±0.07 0.1225±0.04
OCUCFormer[59] 27.30±2.05 0.6998±0.05 0.0289±0.01 0.1988±0.04 29.98±2.23 0.8072±0.05 0.0161±0.03 0.1198±0.03
ResNet[56] 26.60±2.09 0.6611±0.10 0.0305±0.02 0.2097±0.06 27.91±2.04 0.7600±0.07 0.0227±0.01 0.1176±0.04
UNet[26] 28.10±2.47 0.6875±0.11 0.0236±0.02 0.2058±0.06 30.20±2.71 0.7869±0.07 0.0155±0.01 0.1094±0.04
SwinMR[36] 27.42±2.55 0.6813±0.11 0.0278±0.02 0.1994±0.05 29.54±2.68 0.7802±0.07 0.0179±0.01 0.0988±0.03
RefineGAN[57] 28.29±2.53 0.7766±0.06 0.0211±0.01 0.1533±0.03 31.52±2.41 0.8253±0.03 0.0150±0.01 0.0681±0.02
SwinGAN[58] 29.22±2.34 0.7455±0.10 0.0200±0.01 0.1502±0.04 31.89±2.44 0.8277±0.04 0.0144±0.02 0.0677±0.02
OCUCFormer[59] 29.70±2.55 0.7489±0.06 0.0198±0.02 0.1500±0.03 32.00±2.23 0.8288±0.03 0.0166±0.01 0.0676±0.02
TABLE III: Quantitative results of the comparison experiments on the IXI test dataset using different masks. A higher value of metric↑ indicates better performance, whereas a lower value of metric↓ corresponds to improved performance.
Method PSNR↑ SSIM↑ NMSE↓ LPIPS↓ PSNR↑ SSIM↑ NMSE↓ LPIPS↓
ResNet[56] 29.23±3.18 0.9249±0.04 0.0323±0.01 0.0416±0.02 32.69±3.27 0.9609±0.02 0.0148±0.01 0.0229±0.01
UNet[26] 30.54±3.43 0.9434±0.03 0.0244±0.01 0.0291±0.01 33.72±3.53 0.9685±0.02 0.0119±0.01 0.0159±0.01
SwinMR[36] 34.15±3.51 0.9652±0.02 0.0123±0.01 0.0133±0.01 37.70±3.57 0.9805±0.01 0.0055±0.00 0.0055±0.01
RefineGAN[57] 32.96±2.58 0.9500±0.01 0.0126±0.01 0.0142±0.01 40.33±3.22 0.9788±0.02 0.0039±0.01 0.0028±0.01
SwinGAN[58] 33.53±3.32 0.9633±0.02 0.0120±0.01 0.0145±0.01 41.42±3.43 0.9891±0.01 0.0031±0.00 0.0022±0.00
OCUCFormer[59] 33.97±3.54 0.9676±0.02 0.0119±0.01 0.0138±0.02 41.26±3.88 0.9897±0.01 0.0028±0.01 0.0020±0.01
ResNet[56] 31.47±3.65 0.9446±0.03 0.0199±0.01 0.0326±0.02 34.98±3.58 0.9715±0.02 0.0088±0.00 0.0183±0.02
UNet[26] 33.19±4.07 0.9608±0.03 0.0141±0.01 0.0234±0.02 37.12±4.01 0.9809±0.01 0.0057±0.00 0.0135±0.01
SwinMR[36] 32.89±3.42 0.9420±0.02 0.0155±0.01 0.0172±0.01 35.73±3.28 0.9590±0.02 0.0080±0.00 0.0085±0.01
RefineGAN[57] 36.36±4.01 0.9703±0.02 0.0109±0.01 0.0073±0.01 44.80±3.91 0.9910±0.01 0.0017±0.00 0.0012±0.00
SwinGAN[58] 37.44±3.34 0.9800±0.02 0.0056±0.01 0.0067±0.01 45.67±4.43 0.9954±0.00 0.0009±0.00 0.0010±0.00
OCUCFormer[59] 39.48±3.22 0.9860±0.02 0.0034±0.01 0.0065±0.02 46.78±4.40 0.9877±0.02 0.0010±0.01 0.0009±0.01
Figure 4: Visual comparison of different methods under different masks. Red boxes illustrate the enlarged views in detail. Green ellipses and arrows highlight the performance differences in the reconstruction of anatomical structures among various methods.
@@ III-B3 High Resolution Implicit Transformer (HRIT)
III-B4 Tri-Attention Refinement Module (TARM)
@@
As shown in Fig.2 (d), our TARM further refines the reconstruction result $\widehat{{I}}_{3}$ in the image domain, providing it with richer image details.
@@
Fig.3 (b) illustrates the detailed structure of the TARM, which comprises a convolution head and three attention branches, i.e., spatial attention (SA), channel attention (CA), and pixel attention (PA). Specifically, for an image that needs to be refined, i.e., $\widehat{{I}}_{3}$, the TARM firstly extracts shallow feature $f$ by a convolution. Then, $f$ is sent into three attention branches. To detail, PA generates a 3D attention mask $M_{1}$ without any pooling or sampling, which means the output features have local information. CA uses global average pooling (GAP) to generate the 1D attention mask, i.e., $M_{2}$, to obtain features that have global information. SA uses GAP or max average pooling (MAP) to generate a 2D attention mask, i.e., $M_{3}$, to get features with global information. Finally, the weighted features from three attention branches are fused via element-wise summation, to obtain the final reconstruction results $\widehat{{I}}_{4}$ and its corresponding k-space $\widehat{{K}}_{4}$:
$\widehat{{I}}_{4}=conv(M_{1}\otimes f+M_{2}\otimes f+M_{3}\otimes f)\vspace{-0.06cm}$ (15)
$\widehat{{K}}_{4}={FFT}(\widehat{{I}}_{4})$ (16)
III-B5 Loss Functions and Training Strategy

### v2

@@ 3.2 Network Architecture
3.2.1 Low Resolution Implicit Transformer (LRIT)
@@
As depicted in Figure.2 (a), LRIT is an encoder-decoder architecture. It aims to recover LR k-space $\widehat{{K}}_{1}$ and its corresponding $\widehat{{I}}_{1}$ by querying the LR coordinates.
@@
Specifically, given a set of sampled k-values $K_{s}$ and its coordinates $C_{s}$, a tokenization procedure is firstly applied to convert them into a vector sequence $\mathcal{V}$ as follows:
$\mathcal{V}=EM(K_{s})+PE(C_{s}),$ (6)
@@
where $EM$ is the embedding operation via multilayer perceptrons (MLPs) layer, and $PE$ refers to positional encodings with sine and cosine functions. Then we send $\mathcal{V}$ into the encoder of LRIT, which consists of multi-head self-attention ($MSA$) and feed-forward network ($FFN$) as shown in Figure.3 (a). There by, we can obtain the latent space $\mathcal{S}$ as:
$\mathcal{S}=FFN(MSA(\mathcal{V})).$ (7)
@@
Subsequently, we design a decoder that employs standard transformer decoder layers consisting of multi-head cross attention (MCA), MSA, and FFN as shown in Figure.3 (b), to recover complete while low-resolution k-space, i.e., $\widehat{{K}}_{1}$. In detail, for MCA, we employ encoded LR coordinates ${C}_{lr}\in R^{\frac{H}{2}\times\frac{W}{2}}$ to serve as the query ($Q$) and apply two different linear projects to latent space $\mathcal{S}$ to serve as the key ($K$) and value ($V$). After that, we apply MSA to the weighted feature again and decode it via FFN. Thus, we obtain a low-resolution while complete k-space $\widehat{{K}}_{1}\in R^{2\times\frac{H}{2}\times\frac{W}{2}}$, as well as its corresponding MRI image $\widehat{{I}}_{1}\in R^{1\times\frac{H}{2}\times\frac{W}{2}}$:
$\widehat{{K}}_{1}=FFN(MSA(MCA(Q,K,V)),$ (8)
$\widehat{{I}}_{1}={IFFT}(\widehat{{K}}_{1}).$ (9)
@@
3.2.2 Image Domain Guidance Module (IDGM)
As shown in Figure.2 (b), our IDGM fuses the low-quality image $I_{s}$ and the output of LRIT, i.e., $\widehat{{I}}_{1}$, to fully utilize the image information and guide the k-space recovery.
Figure.4 (a) illustrates the detailed structure of the IDGM, which can be divided into two stages. Specifically, during the shallow stage, we firstly perform two convolutions with different strides to extract shallow features from $I_{s}$ and $\widehat{{I}}_{1}$ and unify them to the same size, i.e., ${\widehat{I}_{1}}\in R^{1\times\frac{H}{2}\times\frac{W}{2}}\to R^{h\times\frac{H}{2}\times\frac{W}{2}}$, ${I_{s}}\in R^{1\times H\times W}\to R^{h\times\frac{H}{2}\times\frac{W}{2}}$, where h denotes the channel number. Then, the features are fused via element-wise summation and restored to ${{I}}_{up}\in R^{1\times H\times W}$ via convolution $conv(\cdot)$ and upsampling $up(\cdot)$ as follows:
$I_{up}=up(conv(conv({\widehat{{I}}_{1}})+conv(I_{s}))).$ (10)
During the deep fusion stage, we first employ two convolution heads to extract image-specific shallow features from $I_{s}$ and ${{I}}_{up}$, respectively. After element-wise summation, the fused feature is sent into two attention branches, generating local channel attention $A_{l}$ without any pooling and global channel attention $A_{g}$ by using global average pooling, respectively. Given $A_{l}$ and $A_{g}$, the final attention weights $A$ can be obtained by the broadcasted addition (+), i.e., $A=sigmoid(A_{l}\textbf{+}A_{g})$. Finally, we obtain $\widehat{{I}}_{2}$, as well as its corresponding k-space $\widehat{{K}}_{2}$ as follows:
$\widehat{{I}}_{2}=conv(A\otimes conv({I}_{up})),$ (11)
$\widehat{{K}}_{2}={FFT}(\widehat{{I}}_{2}).$ (12)
3.2.3 High Resolution Implicit Transformer (HRIT)
As shown in Figure.2 (c), to further reconstruct the dense HR k-space, we apply the implicit transformer again. Different from LRIT, here we use the recovered dense $\widehat{{K}}_{2}$ to generate $K_{s}^{\prime}$ and $C_{s}^{\prime}$, forming vector sequence $\mathcal{V^{\prime}}$:
$\mathcal{V^{\prime}}=EM(K_{s}^{\prime})+PE(C_{s}^{\prime}).$ (13)
Then we get latent space $\mathcal{S^{\prime}}$ via the Encoder in Figure.4 (a):
$\mathcal{S^{\prime}}=FFN(MSA(\mathcal{V^{\prime}})).$ (14)
Finally, we encoded HR coordinates $C_{hr}\in R^{H\times W}$ as query ($Q^{\prime}$) and projecting $\mathcal{S^{\prime}}$ as key ($K^{\prime}$) and value ($V^{\prime}$). By using the Decoder shown in Figure.4 (b), we can obtain the high quality k-space $\widehat{{K}}_{3}$ and its corresponding MRI image $\widehat{{I}}_{3}$:
$\widehat{{K}}_{3}=FFN(MSA(MCA(Q^{\prime},K^{\prime},V^{\prime})),$ (15)
$\widehat{{I}}_{3}={IFFT}(\widehat{{K}}_{3}).$ (16)
Figure 5: Visual comparison of different methods under various undersampling masks. Red boxes illustrate the enlarged views in detail. Yellow ellipses highlight the performance differences in the results of anatomical structures among various methods.
3.2.4 Tri-Attention Refinement Module (TARM)
@@
As shown in Figure.2 (d), our TARM refines the reconstructed $\widehat{{I}}_{3}$ in image domain, providing it with richer details.
@@
Figure.4 (b) illustrates the structure of the TARM, which comprises a convolution head and three attention branches, i.e., spatial attention (SA), channel attention (CA), and pixel attention (PA). Specifically, for an image that needs to be refined, i.e., $\widehat{{I}}_{3}$, the TARM first extracts shallow feature $f$ by a convolution. Then, $f$ is sent into three attention branches. To detail, PA generates a 3D attention mask $M_{1}$ without any pooling or sampling, which means the output features have local information. CA uses global average pooling (GAP) to generate the 1D attention mask, i.e., $M_{2}$, to obtain features that have global information. SA generates a 2D attention mask, i.e., $M_{3}$, to get features with global information. Finally, the weighted features from three attention branches are fused via element-wise summation, to obtain the final results $\widehat{{I}}_{4}$ and its corresponding k-space $\widehat{{K}}_{4}$ as follows:
$\widehat{{I}}_{4}=conv(M_{1}\otimes f+M_{2}\otimes f+M_{3}\otimes f),$ (17)
$\widehat{{K}}_{4}={FFT}(\widehat{{I}}_{4}).$ (18)
Table 1: Quantitative results of the comparison experiments on the CC359 validation dataset using different masks. A higher value of metric↑ indicates better performance, whereas a lower value of metric↓ corresponds to improved performance.
Method PSNR↑ SSIM↑ NMSE↓ LPIPS↓ PSNR↑ SSIM↑ NMSE↓ LPIPS↓
ResNet 28.13±3.84 0.8559±0.14 0.0200±0.01 0.0858±0.07 30.05±3.58 0.9052±0.07 0.0131±0.01 0.0569±0.04
UNet 29.34±4.59 0.8756±0.11 0.0156±0.01 0.0618±0.05 31.73±4.70 0.9189±0.07 0.0090±0.00 0.0408±0.04
SwinMR 31.11±3.87 0.9006±0.09 0.0124±0.00 0.0413±0.03 35.67±4.13 0.9543±0.04 0.0045±0.00 0.0184±0.02
RefineGAN 29.96±4.59 0.8888±0.07 0.0303±0.02 0.0395±0.04 38.57±5.16 0.9656±0.02 0.0058±0.01 0.0077±0.01
SwinGAN 31.32±3.20 0.9017±0.06 0.0087±0.01 0.0390±0.02 38.67±4.56 0.9688±0.02 0.0043±0.01 0.0156±0.02
ReconFormer 32.79±2.38 0.9189±0.05 0.0072±0.02 0.0373±0.03 38.75±4.77 0.9696±0.02 0.0032±0.03 0.0149±0.02
ResNet 28.83±3.47 0.8576±0.12 0.0185±0.02 0.0841±0.08 31.00±3.66 0.9102±0.09 0.0181±0.05 0.0592±0.10
UNet 30.39±4.78 0.8790±0.12 0.0124±0.00 0.0578±0.05 33.39±4.98 0.9280±0.08 0.0063±0.00 0.0362±0.03
SwinMR 31.44±4.01 0.8943±0.09 0.0116±0.04 0.0441±0.03 34.36±3.85 0.9378±0.05 0.0059±0.00 0.0220±0.01
RefineGAN 30.08±4.61 0.8770±0.07 0.0303±0.02 0.0395±0.03 35.73±5.05 0.9486±0.03 0.0098±0.01 0.0140±0.01
SwinGAN 34.56±3.56 0.9233±0.04 0.0078±0.01 0.0221±0.02 38.89±4.32 0.9676±0.04 0.0046±0.02 0.0089±0.01
ReconFormer 36.01±3.35 0.9400±0.04 0.0045±0.02 0.0192±0.01 41.35±2.78 0.9784±0.04 0.0022±0.01 0.0055±0.01
Table 2: Quantitative results of the comparison experiments on the fastMRI validation dataset using different masks. A higher value of metric↑ indicates better performance, whereas a lower value of metric↓ corresponds to improved performance.
Method PSNR↑ SSIM↑ NMSE↓ LPIPS↓ PSNR↑ SSIM↑ NMSE↓ LPIPS↓
ResNet 24.54±2.02 0.6062±0.09 0.0461±0.02 0.2469±0.08 25.75±2.04 0.7059±0.07 0.0351±0.02 0.1791±0.06
UNet 26.53±2.22 0.6599±0.10 0.0316±0.02 0.2242±0.08 28.21±2.33 0.7559±0.07 0.0220±0.02 0.1541±0.06
SwinMR 24.67±2.21 0.6214±0.10 0.0480±0.03 0.2166±0.06 27.33±2.46 0.7530±0.08 0.0271±0.02 0.1354±0.05
RefineGAN 25.71±2.56 0.6961±0.06 0.0335±0.02 0.1922±0.04 29.38±2.37 0.8054±0.02 0.0174±0.08 0.1376±0.01
SwinGAN 26.60±2.09 0.7001±0.08 0.0330±0.01 0.1996±0.03 29.89±2.33 0.8069±0.06 0.0169±0.07 0.1225±0.04
ReconFormer 27.65±2.34 0.6999±0.04 0.0269±0.01 0.1887±0.04 30.03±2.54 0.8072±0.03 0.0154±0.03 0.1170±0.03
ResNet 26.60±2.09 0.6611±0.10 0.0305±0.02 0.2097±0.06 27.91±2.04 0.7600±0.07 0.0227±0.01 0.1176±0.04
UNet 28.10±2.47 0.6875±0.11 0.0236±0.02 0.2058±0.06 30.20±2.71 0.7869±0.07 0.0155±0.01 0.1094±0.04
SwinMR 27.42±2.55 0.6813±0.11 0.0278±0.02 0.1994±0.05 29.54±2.68 0.7802±0.07 0.0179±0.01 0.0988±0.03
RefineGAN 28.29±2.53 0.7766±0.06 0.0211±0.01 0.1533±0.03 31.52±2.41 0.8253±0.03 0.0150±0.01 0.0681±0.02
SwinGAN 29.22±2.34 0.7455±0.10 0.0200±0.01 0.1502±0.04 31.89±2.44 0.8277±0.04 0.0144±0.02 0.0677±0.02
ReconFormer 29.73±2.62 0.7485±0.04 0.0199±0.02 0.1514±0.02 32.02±1.89 0.8292±0.04 0.0156±0.02 0.0681±0.02
Table 3: Quantitative results of the comparison experiments on the IXI test dataset using different masks. A higher value of metric↑ indicates better performance, whereas a lower value of metric↓ corresponds to improved performance.
Method PSNR↑ SSIM↑ NMSE↓ LPIPS↓ PSNR↑ SSIM↑ NMSE↓ LPIPS↓
ResNet 29.23±3.18 0.9249±0.04 0.0323±0.01 0.0416±0.02 32.69±3.27 0.9609±0.02 0.0148±0.01 0.0229±0.01
UNet 30.54±3.43 0.9434±0.03 0.0244±0.01 0.0291±0.01 33.72±3.53 0.9685±0.02 0.0119±0.01 0.0159±0.01
SwinMR 34.15±3.51 0.9652±0.02 0.0123±0.01 0.0133±0.01 37.70±3.57 0.9805±0.01 0.0055±0.00 0.0055±0.01
RefineGAN 32.96±2.58 0.9500±0.01 0.0126±0.01 0.0142±0.01 40.33±3.22 0.9788±0.02 0.0039±0.01 0.0028±0.01
SwinGAN 33.53±3.32 0.9633±0.02 0.0120±0.01 0.0145±0.01 41.42±3.43 0.9891±0.01 0.0031±0.00 0.0022±0.00
ReconFormer 33.89±2.87 0.9658±0.02 0.0120±0.02 0.0139±0.01 41.19±2.76 0.9876±0.02 0.0027±0.01 0.0022±0.01
ResNet 31.47±3.65 0.9446±0.03 0.0199±0.01 0.0326±0.02 34.98±3.58 0.9715±0.02 0.0088±0.00 0.0183±0.02
UNet 33.19±4.07 0.9608±0.03 0.0141±0.01 0.0234±0.02 37.12±4.01 0.9809±0.01 0.0057±0.00 0.0135±0.01
SwinMR 32.89±3.42 0.9420±0.02 0.0155±0.01 0.0172±0.01 35.73±3.28 0.9590±0.02 0.0080±0.00 0.0085±0.01
RefineGAN 36.36±4.01 0.9703±0.02 0.0109±0.01 0.0073±0.01 44.80±3.91 0.9910±0.01 0.0017±0.00 0.0012±0.00
SwinGAN 37.44±3.34 0.9800±0.02 0.0056±0.01 0.0067±0.01 45.67±4.43 0.9954±0.00 0.0009±0.00 0.0010±0.00
ReconFormer 39.52±2.76 0.9857±0.03 0.0033±0.02 0.0064±0.02 46.65±3.23 0.9868±0.01 0.0010±0.02 0.0009±0.02
Table 4: Performance comparison of MRI reconstruction under $25\%$ and $12.5\%$ 1D Cartesian mask on the multi-coil SKM-TEA.
sampling ratio = 25% sampling ratio = 12.5%
Methods PSNR↑ SSIM↑ NMSE↓ PSNR↑ SSIM↑ NMSE↓
UNet 33.91 0.8469 0.0204 31.44 0.7904 0.0204
KIKI-Net 34.26 0.8577 0.0196 31.42 0.7941 0.0196
SwinMR 34.45 0.8597 0.0192 31.94 0.8022 0.0192
D5C5 34.63 0.8648 0.0188 31.89 0.8030 0.0188
ReconFormer 35.06 0.8730 0.0179 32.51 0.8158 0.0179
Ours 35.14 0.8832 0.0167 32.66 0.8198 0.0173
3.2.5 Loss Functions and Training Strategy